One number is 9 more than the another number. the smaller number divided by the larger number equals 2/5. find the values of the two numbers?
Answer by mananth(16949)

let one number be x
the other number = x+9
..
x /x+9 = 2/5
5x=2(x+9)
5x=2x+18
3x=18
x=6
the other number = x+9= 15